### v2

# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Inference Time Adaptation

Abstract
Diffusion-based Neural Combinatorial Optimization (NCO) has demonstrated effectiveness in solving NP-complete (NPC) problems by learning discrete diffusion models for solution generation, eliminating hand-crafted domain knowledge. Despite their success, existing NCO methods face significant challenges in both cross-scale and cross-problem generalization, and high training costs compared to traditional solvers. While recent studies on diffusion models have introduced training-free guidance approaches that leverage pre-defined guidance functions for conditional generation, such methodologies have not been extensively explored in combinatorial optimization. To bridge this gap, we propose a training-free inference time adaptation framework (DIFU-Ada) that enables both the zero-shot cross-problem transfer and cross-scale generalization capabilities of diffusion-based NCO solvers without requiring additional training. We provide theoretical analysis that helps understanding the cross-problem transfer capability. Our experimental results demonstrate that a diffusion solver, trained exclusively on the Traveling Salesman Problem (TSP), can achieve competitive zero-shot transfer performance across different problem scales on TSP variants, such as Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), through inference time adaptation.
1 Introduction
Combinatorial optimization (CO) problems are fundamental challenges across numerous domains, from logistics and supply chain management to network design and resource allocation. While traditional exact solvers and heuristic methods have been widely studied, they often struggle with scalability and require significant domain expertise to design problem-specific algorithms [1, 2].
Recent advances in deep learning have sparked interest in Neural Combinatorial Optimization (NCO), which aims to learn reusable solving strategies directly from data, eliminating the need for hand-crafted heuristics [3]. Among various deep learning approaches, diffusion-based models [4, 5] have emerged as a particularly promising direction for solving combinatorial optimization problems. These models have demonstrated remarkable capabilities in learning complex solution distributions by adapting discrete diffusion processes to graph structures [6]. Recent works like [7, 8] have achieved state-of-the-art performance on classical problems such as the Traveling Salesman Problem (TSP), showcasing the potential of diffusion-based generative approaches in combinatorial optimization.
However, the practical applicability of existing NCO approaches is limited by several generalization challenges. First, current models suffer from cross-scale generalization, with performance degrading significantly when applied to larger problem instances than those seen during training, especially for auto-regression-based solvers [9, 10] including transformer and reinforcement learning methods. Second, these models show limited cross-problem transfer capabilities, struggling to adapt to problem variants with modified objectives or additional constraints. While several studies have attempted to enhance learning-based solvers’ generalization through approaches such as training additional networks [11] and fine-tuning [12], these methods require substantial computational costs and training data for training separate models for each problem type and scale.
In parallel, recent advances in diffusion models, particularly in computer vision, have demonstrated the effectiveness of training-free guidance approaches for enhancing conditional generation [13, 14, 15, 16, 17]. These approaches leverage plug-and-play guidance functions or pre-trained networks to enable conditional generation without additional training overhead. These approaches demonstrate significant potential for modifying the inference phase to achieve controllable sampling from pre-trained diffusion models. Inspired by these developments, we explore the training-free inference time adaptation to address the cross-problem transfer challenges in neural combinatorial optimization in Figure 1.
Figure 1: Overview of Inference Time Adaptation for achieving zero-shot cross-problem generalization, shifting from pre-trained problem $G$ (TSP) solution space to target problem $G^{\prime}$ (PCTSP).
Figure 2: The proposed Inference Time Adaptation framework. This approach combines (1) energy-guided sampling, which incorporates problem-specific objectives and constraints, with (2) a recursive denoising-decoding travel strategy for solution refinement, enabling zero-shot cross-problem transfer without training. The Optimality Gap ($\downarrow$) on PCTSP-20 is reduced from $19.21\%$ to $4.20\%$.
This work introduces an inference time adaptation framework (illustrated in Figure 2) designed to enhance the generalization capabilities of diffusion-based Neural Combinatorial Optimization (NCO) solvers without additional training costs. We scale up the inference time by combining two key components: (1) energy-guided sampling, which integrates problem-specific objective functions and constraints into the denoising process, and (2) a recursive denoising-decoding strategy applied at intermediate noise levels. This inference time adaptation framework enables zero-shot cross-problem transfer while maintaining solution feasibility. Furthermore, we present a theoretical analysis to deepen the understanding of the problem transfer capability afforded by this inference time adaptation when applied to pre-trained models. Through experimental evaluation of the TSP-trained diffusion model on more complex variants, the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), we empirically demonstrate its effective zero-shot transferability across different problems of increasing complexity, while maintaining consistent performance across different problem scales. Our work represents a significant step toward more flexible and generalizable diffusion-based combinatorial optimization solvers, reducing the need for problem-specific model training while maintaining competitive performance.
2 Related Works
Neural Network-based Combinatorial Solvers. Neural Combinatorial Optimization (NCO) approaches focus on leveraging neural networks to learn feasible solution distributions for combinatorial optimization problems [3, 18]. Autoregressive construction solvers [9, 10, 19, 20, 21] are built upon the success of transformer-based [22] and reinforcement learning architectures in sequential generation tasks. However, non-autoregressive construction solvers [23, 24, 25, 11, 6, 26] have also been proposed to learn high-quality solution distributions.
Diffusion-based Generative Modeling. Recent advances in generative modeling have revolutionized various domains through diverse approaches, including Variational Autoencoders (VAE) [27], Generative Adversarial Networks (GAN) [28], Diffusion models [4], and GFlowNet [29]. In particular, score-based diffusion models [4, 5, 30, 31, 32, 33] have emerged as a powerful framework operating in continuous domains.
Beyond their state-of-the-art performance in traditional generative tasks, these models have shown remarkable potential in combinatorial optimization (CO). Pioneering work by [6] established new state-of-the-art results for the Traveling Salesman Problem (TSP) by adapting discrete diffusion models [34] to graph structures. Building upon this foundation, [7] enhanced the framework’s performance through gradient search and the current state-of-the-art [8] introduced consistency models to achieve single-step generation. [26] proposed the first diffusion-based unsupervised learning framework.
Training-free Guidance for Diffusion Models. Conditional generation has emerged as a crucial component in real-world applications, enabling precise control over generated outputs. While traditional approaches such as classifier guidance [32] and classifier-free guidance [35] have proven effective, they require substantial computational overhead due to additional training requirements for either the classifier or the diffusion model.
A promising alternative has emerged through training-free guidance methods [13, 14, 15, 16], which are guided by pre-trained networks or loss functions. In the context of discrete diffusion models, this approach remained largely unexplored until [7] pioneered the adaptation of loss-based guidance during inference, building upon the pre-trained discrete diffusion solver framework [6]. Our work demonstrates the significant potential of modifying the infernce phase in enhancing the zero-shot cross-problem transfer of diffusion-based NCO solvers.
3 Preliminaries
3.1 Graph-based CO Problems
Combinatorial optimization (CO) problems on graphs are fundamental to numerous real-world applications. Following recent advances [6, 7], we address these problems by formalizing graph-based CO instances. Following [6, 7, 8], we represent each problem instance as an undirected graph $G(V,E)\in\mathcal{G}$, where $V$ and $E$ denote the vertex and edge sets, respectively. This representation encompasses both vertex selection and edge selection problems, covering a broad spectrum of practical CO scenarios. For any instance $G\in\mathcal{G}$, we define a binary decision variable $\mathbf{x}\in\mathcal{X}_{\mathcal{G}}$, where $\mathcal{X}_{\mathcal{G}}=\{0,1\}^{N}$ represents the feasible solution space. The optimization objective is to find the optimal solution $\mathbf{x}^{*}$ that minimizes a problem-specific objective function $\phi(\cdot;G):\{0,1\}^{N}\rightarrow\mathbb{R}$:
$\mathbf{x}^{*}=\underset{\mathbf{x}\in\mathcal{X}_{\mathcal{G}}}{\text{argmin}}\;\phi(\mathbf{x};G),$ (1)
where the objective function decomposes into a log-barrier formulation [36]:
$\phi(\mathbf{x};G)=f_{\text{cost}}(\mathbf{x};G)-\mu\sum_{i=1}^{m}\log\big{(}-g_{i}(\mathbf{x};G)\big{)},$ (2)
Here $f_{\text{cost}}(\cdot;G)$ measures the solution quality, and $g_{i}(\cdot;G)$ enforces problem-specific constraints through a penalty coefficient $\mu>0$. The validity functions $g_{i}(\cdot;G)$ returns 00 for feasible solutions and is strictly positive for infeasible ones.
3.2 Probabilistic Modeling for CO Problems
To leverage recent advances in deep generative models, we reformulate the CO objective through an energy-based perspective [37]. Specifically, we establish an energy function:
$\mathcal{E}(\cdot;G):=\rho(y,\phi(\cdot;G)),$ (3)
where $\rho$ represents any loss functions (cross entropy loss, L2-norm), and the energy function maps each solution to its corresponding energy state. This energy-based formulation naturally leads to a probabilistic framework through the Boltzmann distribution [38]:
$p(y|\mathbf{x};G)=\frac{\exp\big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\big{)}}{\mathcal{Z}},\quad\text{where}\;\mathcal{Z}=\sum_{\mathbf{x}}\exp\Big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\Big{)},$ (4)
where $\tau$ controls the temperature of the system and $\mathcal{Z}$ denotes the partition function that normalizes the distribution. Recent works have demonstrated promising approaches to approximate this distribution using diffusion-based deep generative models by parameterizing a conditional distribution $p_{\theta}(\mathbf{x}|G)$ to minimize the energy function. Both supervised [6, 7, 8] and unsupervised [26] learning paradigms have shown significant advances. Since our proposed training-free guidance mechanism is applicable to any pre-trained diffusion-based solver, we focus on the supervised learning framework in this work.
Given a training set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of i.i.d. problem instances with their optimal solutions $\mathbf{x}$ and the corresponding optimal objective values $y_{G}^{*}$, we use generative model and optimize the model parameters $\theta$ by maximizing the likelihood of the optimal solutions:
$L(\theta)=\mathbb{E}_{G\sim\mathcal{G}}[-\log p_{\theta}(\mathbf{x}|y_{G}^{*},G)].$ (5)
3.3 Discrete Diffusion Generative Models
We adopt a discrete diffusion framework [34] to effectively sample optimal solutions from the learned distribution $p_{\theta}(\mathbf{x}|y^{*},G)$. In contrast to continuous diffusion models that employ Gaussian noise, the discrete formulation is particularly well-suited for graph-based combinatorial optimization problems [6, 7, 8].
The diffusion process consists of two key components: a forward process that gradually corrupts the data, and a reverse process that learns to reconstruct the original distribution. The forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ maps clean data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0}|G)$ to a sequence of increasingly corrupted latent variables $\mathbf{x}_{1:T}$. The reverse process $p_{\theta}(\mathbf{x}_{0:T}|G)=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)$ learns to gradually denoise these latent variables to recover the original distribution. From a variational perspective, we optimize the model by minimizing an upper bound on the negative log-likelihood, where $C$ is a constant:
$L(\theta)$ $=\mathbb{E}_{G\sim\mathcal{G}}[-\log p_{\theta}(\mathbf{x}_{0}|G)]$ (6)
$\leq\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}D_{KL}[q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)]-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},G)\bigg{]}+C.$
For discrete state spaces, we define the forward process using a categorical distribution:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\text{Cat}(\mathbf{x}_{t};\mathbf{p}=\mathbf{\widetilde{x}}_{t-1}\mathbf{Q}_{t}),$ (7)
where $\mathbf{\widetilde{x}}_{t}\in\{0,1\}^{N\times 2}$ represents the one-hot encoding of $\mathbf{x}_{t}\in\{0,1\}^{N}$. The forward transition matrix $\mathbf{Q}_{t}$ is defined as:
$\mathbf{Q}_{t}=\begin{bmatrix}(1-\beta_{t})&\beta_{t}\\ \beta_{t}&(1-\beta_{t})\end{bmatrix},\quad\beta_{t}\in[0,1],$ (8)
where $[\mathbf{Q}_{t}]_{ij}$ denotes the state transition probability from state $i$ to state $j$. The $t$-step marginal distribution and posterior can be derived as:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})$ $=\text{Cat}(\mathbf{x}_{t};\mathbf{p}=\mathbf{\widetilde{x}}_{0}\overline{\mathbf{Q}}_{t}),$ (9)
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ $=\text{Cat}\left(\mathbf{x}_{t-1};\mathbf{p}=\frac{\mathbf{\widetilde{x}}_{t}\mathbf{Q}_{t}^{\top}\odot\mathbf{\widetilde{x}}_{0}\overline{\mathbf{Q}}_{t-1}}{\mathbf{\widetilde{x}}_{0}\overline{\mathbf{Q}}_{t}\mathbf{\widetilde{x}}_{t}^{\top}}\right),$
where $\overline{\mathbf{Q}}_{t}=\mathbf{Q}_{1}\mathbf{Q}_{2}\ldots\mathbf{Q}_{t}$ and $\odot$ denotes element-wise multiplication.
To capture the structural properties of CO problems, we employ an anisotropic graph neural network architecture [23]. For a given instance $G$, the network learns to predict the clean data distribution $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)$. Taking TSP as an example, where $G$ encodes the 2D Euclidean coordinates of vertices, the network outputs an adjacency matrix $p_{\theta}(\mathbf{\widetilde{x}}_{0}|\mathbf{x}_{t},G)\in[0,1]^{N\times 2}$. This matrix parameterizes $N$ independent Bernoulli distributions, each corresponding to a binary decision variable in $\mathbf{\widetilde{x}}_{0}$. The reverse process during sampling follows:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)=\sum_{\widetilde{\mathbf{x}}_{0}}q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\widetilde{\mathbf{x}}_{0})p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t},G).$ (10)
A decoding strategy is essential to construct feasible solutions from the generated adjacency matrix, which is often represented as a heatmap from which an optimal solution is decoded. Among such strategies, greedy decoding is a straightforward approach that iteratively selects nodes or edges with the highest probabilities from this heatmap until a termination condition is met (e.g., forming a complete cycle in Traveling Salesperson Problem (TSP) applications).
4 Proposed Methods: Inference Time Adaptation
4.1 Energy-guided Sampling for Problem Transfer
While training-free guidance has been extensively studied in computer vision [13, 14, 15, 16, 17], its application to combinatorial optimization problems has only recently been explored by [7, 8]. Adopted by [7, 8], we extend this approach by introducing energy-based training-free guidance for zero-shot cross-problem transfer during sampling, enabling flexible incorporation of additional problem-specific objective functions and constraints into pre-trained diffusion-based CO solvers.
Figure 3: Visualization of the recursive denoising-decoding travel strategy. This approach applies re-noising and energy-guided denoising within an efficient intermediate stage (Time Stage 2), based on the observation that the efficacy of energy guidance varies across different noise levels.
Let $\mathcal{G}^{\prime}=\{G^{\prime}_{i}\}_{i=1}^{n}$ denote a set of new problem instances which differs from the pre-trained problem set $\mathcal{G}$. For a new instance $G^{\prime}$ with its optimal solution pair $(\mathbf{x},y_{G^{\prime}}^{*})$, we need to estimate the new reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y_{G^{\prime}}^{*},G^{\prime})$ according to (10). Following the score estimation perspective of diffusion processes [5], we decompose the conditional score function at time step $t$ into two components:
$\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}|y^{*}_{G^{\prime}},G^{\prime})}_{\text{Posterior Score}}$ $=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})}_{\text{Pre-trained Prior Score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Energy-guided Score}}.$ (11)
From the Bayesian perspective, $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ can be understood as the prior, which contains knowledge of the pre-trained problems (i.e. TSP problem in our experimental settings), and $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ corresponds to the adjustive likelihood that incorporates additional constraints or objectives of the variant problem. We sample from the posterior $p_{\theta}(\mathbf{x}_{t}|y^{*}_{G^{\prime}},G^{\prime})$ to generate high-quality solutions to the new problems.
Drawing upon this theoretical framework, we leverage the pre-trained diffusion model to estimate the first term $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})$. In the context of cross-problem transfer, while the pre-trained model yields only a biased score function for new problem instances, we compute the second energy-guided term $\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ to adjust using an energy function that specifically accounts for the additional objectives and constraints of the new problems:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\mathcal{E}(y^{*}_{G^{\prime}},\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$ (12)
$\propto-\nabla_{\mathbf{x}_{t}}\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})),$
where $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}))$ measures the distance of energy states between the optimal value and the predicted solution. Here, $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})$ represents the predicted clean sample from the current noisy state $\mathbf{x}_{t}$, which is estimated by the logits of $N$ independent Bernoulli samples $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t})}[\widetilde{\mathbf{x}}_{0}]$. In this work, we define $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})):=\|y^{*}_{G^{\prime}}-\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})\|_{2}$. From Eq (12), we have:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$ (13)
$p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $=\exp\left(-\frac{1}{\tau}\nabla_{\mathbf{x}_{t}}\phi\left(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}\right)\right).$
where $\tau$ is the energy guidance temperature to control the strength of transferring. Combining Eq. (11) and Eq. (13), we derive an energy-guided reverse sampling process for problem transfer:
$\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})}_{\text{Transfer Posterior}}\propto\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G^{\prime})}_{\text{Pre-trained Prior}}\underbrace{p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Guidance Likelihood}}.$ (14)
In this work, we have the following log-barrier formulations for PCTSP and OP as the guided gradient:
Proposition 4.1 (Prize Collecting TSP (PCTSP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, vertex prizes $r:V\rightarrow\mathbb{R}^{+}$, penalties $p:V\rightarrow\mathbb{R}^{+}$, and prize threshold $R$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
$\phi(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)-\mu\cdot\log(-g(\mathbf{x},\mathbf{y};G))$
$\text{where}\quad f_{\text{cost}}(\mathbf{x},\mathbf{y};G)$ $=\sum_{e\in E}w_{e}x_{e}+\sum_{v\in V}p_{v}(1-y_{v})$
$g(\mathbf{x},\mathbf{y};G)$ $=\max(0,R-\sum_{v\in V}r_{v}y_{v})$
Proposition 4.2 (Orienteering Problem (OP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, vertex scores $s:V\rightarrow\mathbb{R}^{+}$, and budget $B$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
$\phi(\mathbf{x},\mathbf{y};G)$ $=f_{\text{cost}}(\mathbf{x},\mathbf{y};G)-\mu\cdot\log(-g(\mathbf{x},\mathbf{y};G))$
$\text{where}\quad f_{\text{cost}}(\mathbf{x},\mathbf{y};G)$ $=-\sum_{v\in V}s_{v}y_{v}$
$g(\mathbf{x},\mathbf{y};G)$ $=\max(0,\sum_{e\in E}w_{e}x_{e}-B)$
4.2 Recursive Denoising-decoding Travel Strategy
Our initial experiments applying energy-guided denoising for cross-problem transfer highlighted a significant challenge: this approach alone often showed insufficient to effectively bridge the gap between the learned distribution of pre-trained problem types and the target distribution of new problem instances. This distributional divergence frequently hindered the generation of high-quality, feasible solutions for cross-problem transfer.
To address this, we first scaled up the inference time and explored a iterative rewrite strategy, adapting from [7]. This involved taking an initial solution $x_{0}^{(1)}$ for the pre-trained problem instance $G$, adding noise to obtain a highly noisy graph $x_{T}^{(1)}$, and then denoising to obtain the clean state $x_{0}^{(2)}$, aiming to progressively shift the solution towards the characteristics of the target problem. While this recursive strategy demonstrated improved transfer capabilities, it came at the cost of a considerable increase in inference time due to the repeated denoising iterations.
This led us to investigate the efficiency of the guidance mechanism itself. Inspired by the time-travel strategies in vision tasks [15], which revealed that guidance effects are not uniform across diffusion time steps, we hypothesized a similar phenomenon in NCO task. Our empirical analysis confirmed this: the impact of energy guidance varied significantly with the noise level. Crucially, we observed that guidance was most potent and beneficial at intermediate noise levels from 200 to 500 steps. In contrast, its effect was minimal on highly noisy graphs (analogous to an early time stage) and became largely redundant for nearly converged solutions (a late, refinement stage), where the solution structure is already well-defined.
Based on this observation that the impact of energy guidance at different noise levels, we developed our optimized recursive denoising-decoding travel strategy in Figure 3. Instead of applying full re-noising and iterative energy-guided denoising across many steps, our refined approach reduces the re-noising process to only a few steps. More importantly, the computationally intensive energy-guided denoising is applied for just a single step, precisely targeted at these effective intermediate noise levels. This targeted acceleration reduces the computational overhead associated with the recursive process, yielding a 5-10x speedup in inference time compared to the original iterations. This optimization not only maintains but enhances solution quality in several instances, establishing a significantly more favorable performance-to-speed trade-off.
Algorithm 1 outlines our proposed 2-phase Inference Time Adaptation framework, which combines recursive denoising-decoding travel strategy with energy-guided sampling. This approach offers a practical and efficient method for achieving zero-shot cross-problem transfer using pre-trained diffusion models, without any additional training costs. Furthermore, the framework exhibits flexibility by incorporating advanced heuristic decoding strategies, such as 2-opt heuristics [39] or Monte Carlo Tree Search (MCTS) [24, 6], to enhance solution quality.
Algorithm 1 Inference Time adaptation for Cross-problem Transfer
0:
1:  $p_{\theta}$: Pre-trained diffusion model
2:  $G^{\prime}$: Target problem instance
3:  $T$: Number of diffusion steps
4:  $i,j$: Number of intermediate steps
5:  $\tau$: Energy guidance temperature
6:  $K$: Number of recursive travel
6:  Optimal solution $\mathbf{y}_{K}$ for problem instance $G^{\prime}$
7:  Initialize $\mathbf{x}_{T}$ with random binary values
8:  for $k=1$ to $K$ do
9:     Re-noise previous estimated solution $\mathbf{y}_{k-1}$ to $\mathbf{x}_{j}$
10:     for $t=j$ to $i$ do
11:        Compute $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ from pre-trained model
12:        Compute energy gradient: $\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})$
13:        Compute $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})\propto\exp(-\nabla_{\mathbf{x}_{t}}\phi/\tau)$
14:        Compute posterior $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})$ by Eq (14)
15:        Update next state with Bernoulli Sampling: $\mathbf{x}_{t-1}\sim\text{Cat}(\mathbf{x}_{t-1};p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime}))$
16:     end for
17:     Estimate $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{i})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{i})}[\widetilde{\mathbf{x}}_{0}]$
18:     Decode $\widetilde{\mathbf{x}}_{0}$ to the optimal solution $\mathbf{y}_{k}$
19:  end for
20:  return  $\mathbf{y}_{K}$
4.3 Theoretical Analysis of Cross-problem Transfer
Consider the scenario where we aim to transfer knowledge from the original Traveling Salesman Problem (TSP) to its variants: the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP). The diffusion model, trained on optimal TSP instances, estimates $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ such that, under the assumption of perfect optimization, $p_{\theta}(\mathbf{x}_{0}|G^{\prime})$ approximates $q_{\textup{TSP}}(\mathbf{x}_{0}|G^{\prime})$, where the latter represents the distribution of ground-truth TSP solutions for instance $G^{\prime}$. Through the Bayesian equation established in (11), we expect such a pre-trained prior would substantially enhance posterior sampling solution quality, as TSP shares fundamental structural similarities with PCTSP and OP, thereby encoding relevant domain knowledge. We formalize such similarities in the following analysis. The complete proof is provided in Appendix B. We make use of the following definition to facilitate the analysis.
Definition 4.3 (Marginal Decrease).
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ denote the cost of the optimal TSP tour visiting all nodes in $S$. The marginal decrease of a subset $S$ is defined as
$\Delta(S)=\textup{TSP}(V)-\textup{TSP}(V\setminus S).$
The marginal decrease measures the cost reduction of not visiting a subset of nodes $S$, which helps quantify the difference between the optimal tours. Take PCTSP as an example. If for any non-empty subset of nodes $S\subseteq V$, the penalty of not visiting the nodes in $S$ satisfies $\sum_{i\in S}{p_{i}}\geq\Delta(S)$, then PCTSP and TSP share the same optimal tours. Based on this notion, we formalize the structural similarities in the following theorem.
Theorem 4.4.
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ and $\textup{argTSP}(S)$ denote the optimal cost and optimal tours of TSP on the subgraph specified by $S$. Under Assumptions [^1] We assume simpler setups for PCTSP and OP for clearer presentation of the underlying similarities. These assumptions can be relaxed as discussed in Appendix C. B.2 and B.3, the optimal tours of PCTSP are $\textup{argTSP}(V\setminus S_{\textup{PCTSP}})$, where
$S_{\textup{PCTSP}}\in\arg\min_{S\subseteq V}{\sum_{i\in S}{p_{i}}-\Delta(S)},$
and the optimal tours of OP are $\textup{argTSP}(V\setminus S_{\textup{OP}})$, where
$S_{\textup{OP}}\in$ $\arg\min_{S\subseteq V}{\Delta(S)}$
[…]
Here $D_{\textup{OP}}$ is the distance limit of the total tour length, and $\Delta(S)$ denotes the marginal decrease of a non-empty subset of nodes $S$.
Theorem 4.4 shows that the optimal solutions of PCTSP and OP are indeed the optimal tours of TSP on some subgraph of $G^{\prime}$, revealing their fundamental structural similarities. We expect that the pre-trained diffusion model is powerful enough to generate high-quality TSP solutions given test instances with various graph sizes. We can therefore understand the inference time adaptation in this scenario as forcing the pre-trained model to focus on a subgraph of $G^{\prime}$. Through this theoretical analysis, we deepen our understanding of how the proposed training-free inference time adaptation works when transferring from a pre-trained network to its variants. Although we focus on PCTSP and OP in this paper, we show that our methodology and insight also hold for more challenging TSP variants such as TSP with time windows in Appendix D.
5 Numerical Results
Dataset. We evaluate our approach on the commonly-used NP-complete combinatorial optimization problems: the Traveling Salesman Problem (TSP) together with its variants, the Prize Collecting Traveling Salesman Problem (PCTSP) and Orienteering Problem (OP). The detailed problem descriptions are given in Appendix A.
Evaluation. Following [10], we generate 1000 test instances for each problem scale: 20, 50, and 100, which denotes the node counts for PCTSP and OP. We evaluate model performance using two primary metrics: average solution cost and optimality gap relative to the exact solution solvers. Additionally, we measure computational efficiency through total training time and per-instance inference time.
Baselines. We compare our approach against multiple baseline categories. For PCTSP, (1) Exact solver: Gurobi; (2) OR-based heuristics: OR-Tools and Iterated Local Search (ILS); (3) Learning-based methods: AM [10], MDAM [40], AM-FT [12], ASP [11], DIFUSCO [6], T2T [7]. For OP, (1) Exact solver: Gurobi; (2) OR-based heuristics: Compass [41] and Tsili [42]; (3) Learning-based methods: AM [10], AM-FT [12], DIFUSCO [6], T2T [7].
Experimental Setup. The proposed Inference Time Adaptation for Diffusion-based solvers (DIFU-Ada) builds upon DIFUSCO’s [6] TSP-trained checkpoints in three different scales without additional training. The backbone archetechture used in diffusion model is the same as DIFUSCO, a 12-layer Anisotropic GNN with a width of 256 hidden nodes. The results of DIFU-Ada are recorded under 100 iterations, with 5-step re-noising and 1-step guided denoising. The decoding strategy used in experiments is greedy decoding. The energy functions used in guided sampling are stated in Proposition 4.1 and 4.2. The guidance temperature $\lambda$ is fixed to be 0.1 and the constraint coefficient $\mu$ is set to be 1 for all problem instances. All experiments are conducted on a single Tesla V100 GPU.
PCTSP-20 PCTSP-50 PCTSP-100
[…]
DIFUSCO 3.78 19.21% 1.04s 5.20 15.97% 1.35s 7.85 32.31% 2.02s
T2T 3.64 14.82% 2.02s 4.90 9.38% 2.18s 7.23 21.92% 2.70s
DIFU-Ada (Ours) 3.30 4.20% 1.72s 4.63 3.57% 1.99s 6.51 9.61% 2.74s
OP-20 OP-50 OP-100
[…]
DIFUSCO 9.25 12.48% 1.51s 25.60 13.45% 1.88s 45.66 20.02% 2.04s
T2T 9.67 8.51% 2.70s 26.89 9.09% 3.04s 48.72 14.70% 3.45s
DIFU-Ada (Ours) 10.24 3.11% 2.09s 28.21 4.63% 2.35s 54.56 8.06% 2.91s
Table 1: Zero-shot cross-problem transfer performance comparison between TSP-trained DIFUSCO, T2T and DIFU-Ada (Ours) approach on PCTSP and OP instances. The results show solution cost, optimality gap, and the average inference time for one instance across three problem scales (20, 50, and 100 nodes).
5.1 Cross-problem and Cross-scale Generalization Results
To validate the effectiveness of our inference time adaptation framework in zero-shot cross-problem transfer scenarios, we first conduct experiments while maintaining consistent problem scales. As demonstrated in Table 1, our method achieves significant zero-shot performance improvements when applied to the pre-trained DIFUSCO [6] and T2T [7] model on TSP. These improvements are evident across all PCTSP and OP problem scales, in both enhanced solution quality and reduced optimality gaps compared to the exact OR solvers.
Following the evaluation strategy following [11], we conducted experiments of DIFU-Ada’s transferability across diverse combinatorial optimization (CO) problem types and scales. Our evaluation framework specifically assesses two aspects of generalization: 1) Cross-problem Generalization: We assess the model’s adaptability to different CO problem variants, such as transferring from TSP to PCTSP or OP. 2) Cross-scale Generalization: We evaluate the performance of a single, optimally trained model across various problem scales (e.g., from small to large instances of the same problem type). Unlike existing learning-based methods that necessitate problem-specific retraining for new problem types or scales (e.g., for PCTSP and OP instances), our DIFU-Ada framework achieves zero-shot generation at inference time. This rigorous out-of-distribution evaluation strategy presents a more challenging benchmark than conventional in-distribution assessments, testing the model’s capacity of generalization.
The experimental results presented in Table 2 (PCTSP) and Table 3 (OP) show the cross-problem and cross-scale generalization capabilities of DIFU-Ada. Our framework achieves competitive average optimality gaps ($9.70\%$ for PCTSP and $7.26\%$ for OP) across all scales, while maintaining its zero-shot cross-problem transfer capability. This performance is achieved without the training costs (e.g., 3-5 days of retraining reported for existing methods) or the need for well-labeled training datasets. While OR-based heuristics like Compass exhibit strong performance on narrowly defined problem types, DIFU-Ada’s ability to leverage pre-trained diffusion models across different CO problem variants without additional training highlights its flexibility and broader applicability. Besides, the results on large-scale PCTSP-500 and PCTSP-1000 in Appendix F, Table 4 demonstrate the scalability of DIFU-Ada.
PCTSP-20 PCTSP-50 PCTSP-100
Method Gap $\downarrow$ Time $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Avg Gap $\downarrow$ Training-free Training Time $\downarrow$
OR Gurobi 0.00% 3.10s
OR-Tools 2.13% 12.31s 4.85% 2.02m 10.33% 5.84m 5.77%
ILS (C++) 1.07% 2.13s 0.00% 18.30s 0.00% 56.11s 0.36%
ILS (Python 10x) 63.23% 3.05s 148.05% 4.70s 209.78% 5.27s 140.35%
Learning AM (Greedy) 2.76% 0.02s 18.20% 0.07s 28.98% 0.15s 16.65% 3.5 days
AM (Sampling) 2.54% 2.43s 14.58% 7.08s 22.20% 15.13s 13.11% 3.5 days
MDAM (Greedy) 11.76% 41.10s 24.73% 1.31m 30.07% 1.96m 22.19% 4.3 days
MDAM (Beam Search) 5.88% 2.70m 18.81% 4.77m 26.09% 6.97m 16.93% 4.3 days
ASP 12.05% 0.03s 10.34% 0.08s 11.56% 0.18s 11.32% 4.6 days
AM-FT (greedy) 2.11% 0.03s 16.58% 0.07s 29.08% 0.16s 15.92% 4.9 days
AM-FT (Sampling) 1.02% 2.51s 14.11% 8.02s 25.19% 17.21s 13.44% 4.9 days
DIFU-Ada (Ours) 4.45% 1.80s 8.43% 1.92s 16.22% 2.87s 9.70% 0 day
Table 2: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on PCTSP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (OR-Tools, ILS), learning-based models (AM, MDAM, ASP, AM-FT, DIFU-Ada). Performance metrics include optimality gap, inference time, and training time. The results marked with * are reported from [11]. The proposed DIFU-Ada achieves competitive performance while requiring no training.
OP-20 OP-50 OP-100
Method Gap $\downarrow$ Time $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Gap $\downarrow$ Time $\downarrow$ Avg Gap $\downarrow$ Training-free Training Time $\downarrow$
OR Gurobi 0.00% 10.22s
Compass 0.15% 0.27s 0.00% 1.02s 0.00% 6.74s 0.05%
Tsili (Greedy) 16.58% 0.02s 19.22% 0.03s 19.71% 0.03s 18.50%
Tsili (Sampling) 0.85% 0.55s 4.46% 2.45s 8.56% 9.08s 4.62%
Learning AM (Greedy) 5.78% 0.04s 12.10% 0.09s 28.75% 0.19s 15.54% 3.2 days
AM (Sampling) 2.62% 2.87s 8.81% 8.12s 22.17% 14.98s 11.20% 3.2 days
AM-FT (Greedy) 5.50% 0.04s 11.51% 0.08s 23.40% 0.18s 13.47% 4.7 days
AM-FT (Sampling) 1.01% 2.76s 7.19% 8.44s 19.51% 18.15s 9.24% 4.7 days
DIFU-Ada (Ours) 3.44% 2.01s 6.11% 2.29s 12.21% 2.85s 7.26% 0 day
Table 3: Comprehensive evaluation of cross-scale generalization capabilities across different solver categories on OP instances. Comparison includes exact solvers (Gurobi), OR-based heuristics (Compass, Tsili), learning-based models (AM, AM-FT, DIFU-Ada). The results marked with * are reported from [12].
Figure 4: Impact of the number of recursive travel steps in the inference time adaptation on the trade-off between optimality gap (%) and inference time (s) for PCTSP and OP.
5.2 Ablation Studies
The DIFU-Ada framework integrates two key components: energy-guided sampling and recursive denoising-decoding travel strategy. To assess the contribution of each, we conducted ablation studies. Figure 4 quantitatively illustrates the impact of the recursive denoising-decoding strategy, demonstrating a progressive enhancement in cross-problem transfer capabilities as the number of recursive iterations increases. Figure 5 presents a analysis of performance with or without energy-guided sampling during the recursive iterations. The performance gap observed in these studies show the efficacy of both components and highlight their distinct roles in the overall effectiveness to achieve zero-shot cross-problem transfer.
Moreover, Figure 6 shows the effect of the temperature parameter in controlling the guidance strength of adaptation. Higher temperature effectively preserve the structural information in the original problem instances, whereas lower temperatures adapt to the characteristics of new problem instances. This temperature-dependent behavior enables fine-grained control over the balance between maintaining prior knowledge and new problem features. In Appendix F, Figure 7, the results on varying hyperparameter $\mu$ in energy function show that the performances are not sensitive to the choice of $\mu$, thus we choose the optimal $\mu=1$ empirically for all problems.
Figure 5: The Optimality Gap of performing recursive travel strategy with or without energy-guided sampling on PCTSP and OP.
Figure 6: Optimality gap changes with respect to the guided temperature ($-\lambda$ as the x-axis label) on PCTSP and OP.
6 Conclusions
In this work, we introduced an Inference Time Adaptation framework (DIFU-Ada) enabling zero-shot cross-problem transfer for diffusion-based solvers. Through an integration of energy-guided sampling and recursive denoising-decoding travel strategy, our approach successfully transfers pre-trained diffusion models to generate high-quality solutions for TSP variants without additional training. We also provided theoretical analysis in this scenario to deepen our understanding of the problem transfer capability in pre-trained diffusion-based NCO solvers. Numerical experiments on different scales of PCTSP and OP demonstrate that our framework achieves competitive zero-shot performance against existing learning-based methods across various scales.
This work demonstrates the considerable potential of adapting diffusion-based generative NCO solvers to address real-world combinatorial optimization problems, particularly those involving dynamic constraints and objectives. Our approach reduces the computational costs of retraining large-scale neural solvers by providing a flexible, off-the-shelf sampling scheme that enhances both cross-problem and cross-scale generalization capabilities. Nevertheless, several limitations and directions for future research still exist. While our method exhibits promising results on TSP variants, its applicability to a broader spectrum of combinatorial optimization problems beyond routing domains remains to be validated.
Appendix A Problem Descriptions of TSP Variants
  • Traveling Salesman Problem (TSP) requires finding the minimal-length Hamiltonian cycle in a complete graph, where the salesman must visit each city exactly once before returning to the starting point.
  • Prize Collecting TSP (PCTSP) [43] extends the classical TSP by introducing node-specific prizes and penalties. The objective is to optimize a trade-off between minimizing tour length and unvisited node penalties while ensuring collected prizes exceed a predefined threshold. This formulation creates a more complex optimization landscape where node visitation decisions must balance multiple competing factors.
  • The Orienteering Problem (OP), first introduced by [44], is a fundamental combinatorial optimization problem with widespread applications in real-world scenarios. In the OP, each node in the network is associated with a non-negative prize value, and the objective is to determine an optimal tour that begins and ends at a designated depot node. The tour must satisfy two key constraints: maximize the total collected prizes from the visited nodes, and ensure the total tour length does not exceed a predetermined distance limit. This problem effectively captures the trade-off between reward collection and resource constraints.
Appendix B Proof of Theorem 4.4
To analyze the relation between the optimal solutions of TSP and that of PCTSP/OP, we make use of the following definitions.
Definition B.1 (Marginal Decrease).
For a non-empty subset of nodes $S\subseteq V$, let $\textup{TSP}(S)$ denote the cost of the optimal TSP tour visiting all nodes in $S$. The marginal decrease of a subset $S$ is defined as
$\Delta(S)=\textup{TSP}(V)-\textup{TSP}(V\setminus S).$
The marginal decrease measures the cost reduction of not visiting a subset of nodes $S$. This concept helps quantify the relation between the optimal solutions of TSP/PCTSP/OP. We first make the following assumptions on the PCTSP and OP problems in the analysis. Note that we make these assumptions just for the simplicity of presentation.
Assumption B.2 (PCTSP Setting).
We consider PCTSP without the minimum prize constraint.
Assumption B.3 (OP Setting).
For OP, we assume an identical reward at each city.
Note that the above assumptions can be relaxed by considering a node-weighted TSP, which can be transformed into a standard TSP as discussed in Appendix C.
Under these assumptions, we compare the optimal solutions of PCTSP/OP with that of TSP as follows.
(i) For PCTSP, if for any non-empty subset of nodes $S\subseteq V$, the penalty of not visiting the nodes in $S$ satisfies
$\sum_{i\in S}{p_{i}}\geq\Delta(S),$
then PCTSP and TSP share the same optimal solutions, since not visiting any subset of nodes would result in a higher total cost. We can thus formulate the optimal cost (tour length + penalty) of PCTSP as
$\text{PCTSP}(V)=\text{TSP}(V)+\min_{S\subseteq V}{\sum_{i\in S}{p_{i}}-\Delta(S)}.$ (15)
And we have the optimal solutions of PCTSP being the tours in $\text{TSP}(V\setminus S)$ with
$S\in\arg\min_{S\subseteq V}{\sum_{i\in S}{p_{i}}-\Delta(S)},$
completing the proof for PCTSP.
(ii) For OP, since we assume a uniform reward at each city, the optimization objective becomes visiting as much nodes as possible under the distance limit of the total tour length. This objective is identical to minimizing the travel cost while respecting the budget, because the optimal cost of TSP satisfies the following property:
[…]
We can thus formulate the optimal tour length of OP as
$\text{OP}(V)={}$ $\max_{S\subseteq V}{\text{TSP}(V\setminus S)}$
[…]
where $D_{\text{OP}}$ is the distance limit of OP. And we have the optimal solutions of OP being the tours in $\text{TSP}(V\setminus S)$ with
$S\in$ $\arg\min_{S\subseteq V}{\Delta(S)}$
[…]
completing the proof for OP.
Appendix C Relaxing Assumptions B.2 and B.3
Our analysis in Theorem 4.4 can be extended to the cases with minimum prize constraints or non-uniform rewards by considering node-weighted TSP formulations. The key is that the optimal solutions of PCTSP/OP are the optimal solutions of certain node-weighted TSP over a subgraph as shown below.
(i) For the general case of OP, we can reformulate the problem by assigning each node $i$ a score $s_{i}$ equal to its negative reward and setting all edge costs to zero. Meanwhile, we maintain the edge cost function $w_{e}$ solely to represent the budget constraint of OP. This transforms OP into a node-weighted TSP with an additional budget constraint.
(ii) Similarly, for PCTSP, we assign $s_{i}=-p_{i}$ (the negative penalty of node $i$) while preserving the original edge costs. The optimal solution to PCTSP then corresponds to that of a node-weighted TSP with minimum prize constraints.
Since the pre-trained model was trained on standard TSP instances, we can convert node-weighted TSP to a standard one by splitting each node $i$ into two nodes $i_{in}$ and $i_{out}$, and adding a directed edge from $i_{in}$ to $i_{out}$ with weight $s_{i}$. Given this conversion, we could explain the effectiveness of our method as follows: the energy-guided sampling implicitly considers the node-weighted TSP and guides the generation to zoom in a subgraph of the node-weighted graph which respects the additional constraints.
Appendix D Adaptation to TSP with Time Windows (TSP-TW)
TSP with Time Windows (TSP-TW) is a variant of TSP that requires each node $i$ to be visited within a given time range $[e_{i},l_{i}]$.
Our approach can also handle TSP-TW through a time-expanded graph transformation:
For each node $i$ with time window $[e_{i},l_{i}]$, we create multiple time-indexed replicas $(i_{e_{i}},i_{e_{i}+1},...,i_{l_{i}})$. The original graph’s edges are then transferred to this expanded representation according to temporal feasibility - if node $i$ connects to node $j$ in the original graph, we connect $i_{t}$ to $j_{t+1}$ in the expanded graph (assuming unit travel time for simplicity).
This transformation converts the time window constraints into a structural pattern within the expanded graph. The TSP-TW solution corresponds to a standard TSP tour on this expanded graph with one critical constraint: exactly one replica of each original node must be visited, expressed as:
$\sum_{t\in[e_{i},l_{i}]}\sum_{(j,i_{t})\in E}x_{j,i_{t}}=1.$
This constraint can be directly encoded as an energy function in our framework. Our method then proceeds by:
  • Sampling initial TSP solutions on the time-expanded graph
  • Using energy-guided sampling to enforce the single-visit-per-node constraint
  • Converging toward feasible TSP-TW solutions
Through this transformation, TSP-TW effectively becomes a "subgraph TSP" scenario similar to that of PCTSP/OP, demonstrating the flexibility of our energy-guided approach for handling complex constraints without requiring problem-specific training.
Appendix E Energy Formulations for TSP
In our experimental setup, we employ an anisotropic graph neural network architecture enhanced with edge-gating mechanisms [45, 46]. This encoder transforms the input 2D coordinates and TSP variant paths into a binary edge-selection graph. Additional problem-specific information - such as prizes and penalties in PCTSP and prizes in OP - is incorporated into the gradient calculation of the guidance objective function. The constraint coefficients $\mu$ are chosen to be $1$ in experiments.
Proposition E.1 (Traveling Salesman Problem (TSP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, find $\mathbf{x}\in\{0,1\}^{|E|}$ that minimizes:
$\phi(\mathbf{x};G)$ $=f_{\text{cost}}(\mathbf{x};G)-\mu\cdot\log(-g(\mathbf{x};G))$
$\text{where}\quad f_{\text{cost}}(\mathbf{x};G)$ $=\sum_{e\in E}w_{e}x_{e}$
Appendix F Additional Experimental Results
PCTSP-500 PCTSP-1000
[…]
OR Tools* 15.0 1h 24.9 1h
OR Tools (more iterations)* 14.4 16h 20.6 16h
AM* 19.3 14m 34.8 23m
MDAM* 14.8 2.8m 22.2 17m
GLOP-S* 14.3 1.5m 19.8 2.5m
DIFU-Ada (Ours) 14.9 7.2m 23.5 9.9m
Table 4: Comparison of various methods on 128 instances of PCTSP-500 and PCTSP-1000. The data generation and evaluations are conducted following Ye et al. (2024). Results marked with * are taken directly from Ye et al. (2024).
Training Method Testing Dataset
TSP-20 TSP-50 TSP-100
TSP-20 DIFUSCO 3.14, 0.23% 5.75, 0.43% 7.90, 1.81%
T2T 3.14, 0.12% 5.71, 0.19% 7.82, 0.87%
Fast T2T 3.12, 0.03% 5.69, 0.00% 7.76, 0.08%
Ours 3.14, 0.15% 5.71, 0.22% 7.83, 0.95%
TSP-50 DIFUSCO 3.14, 0.25% 5.69, 0.09% 7.87, 1.44%
T2T 3.14, 0.11% 5.69, 0.02% 7.80, 0.55%
Fast T2T 3.12, 0.05% 5.69, 0.00% 7.76, 0.08%
Ours 3.14, 0.14% 5.69, 0.05% 7.81, 0.60%
TSP-100 DIFUSCO 3.16, 0.44% 5.70, 0.25% 7.78, 0.23%
T2T 3.14, 0.21% 5.70, 0.11% 7.77, 0.08%
Fast T2T 3.12, 0.10% 5.69, 0.01% 7.76, 0.01%
Ours 3.14, 0.20% 5.71, 0.25% 7.77, 0.09%
Table 5: Results of various diffusion-based methods [6, 7, 8] on different TSP testing datasets (TSP-20, TSP-50, TSP-100), only considering cross-scale generalization capability.
Figure 7: Impact of the energy function hyperparameter $\mu$ on the optimality gap for PCTSP-50 and OP-50.